If the measure of the supplement of an angle is 10 degrees more than 5 times the measure of the complement, what is the measure of the angle.
Thanks for you help.

Answer by scott8148(5880)

let x=angle

"the measure of the supplement of an angle" __ 180-x

"is" __ =

"10 degrees more than 5 times the measure of the complement" __ 5(90-x)+10

180-x=450-5x+10 __ 4x=280